# Identification of fixations and saccades in eye-tracking data using adaptive threshold-based methods

### v2

@@ Abstract
Properties of ocular fixations and saccades are highly stochastic during many experimental tasks, and their statistics are often used as proxies for various aspects of cognition. Although distinguishing saccades from fixations is not trivial, experimentalists generally use common ad-hoc thresholds in detection algorithms. This neglects inter-task and inter-individual variability in oculomotor dynamics, and potentially biases the resulting statistics. In this article, we introduce and evaluate an adaptive method based on a Markovian approximation of eye-gaze dynamics, using saccades and fixations as states such that the optimal threshold minimizes state transitions. Applying this to three common threshold-based algorithms (velocity, angular velocity, and dispersion), we evaluate the overall accuracy against a multi-threshold benchmark as well as robustness to noise. We find that a velocity threshold achieves the highest baseline accuracy (90-93%) across both free-viewing and visual search tasks. However, velocity-based methods degrade rapidly under noise when thresholds remain fixed, with accuracy falling below 20% at high noise levels. Adaptive threshold optimization via $K$-ratio minimization substantially improves performance under noisy conditions for all algorithms. Adaptive dispersion thresholds demonstrate superior noise robustness, maintaining accuracy above 81% even at extreme noise levels ($\sigma=50$ px), though a precision-recall trade-off emerges that favors fixation detection at the expense of saccade identification. In addition to demonstrating our parsimonious adaptive thresholding method, these findings provide practical guidance for selecting and tuning classification algorithms based on data quality and analytical priorities.
@@ 1 Introduction
Conventional algorithms, such as Velocity-Threshold Identification (I-VT), Angular Velocity-Threshold Identification (I-AVT), and Dispersion-Threshold Identification (I-DT), rely on fixed velocity (e.g., 20–40°/s for I-VT) or spatial dispersion thresholds to classify eye movements [10.1145/355017.355028]. While computationally efficient and suitable for real-time applications, these methods are sensitive to noise and often fail to detect auxiliary events like microsaccades or post-saccadic oscillations (glissades), limiting their robustness in complex datasets [martinezs, nystrom2010adaptive]. Recent advancements have introduced more sophisticated approaches to address these limitations. For instance, Larsson et al. developed robust segmentation techniques for fixations and smooth pursuits in high-speed eye-tracking data, outperforming velocity- and dispersion-based methods [LARSSON2015145]. Zemblys et al. leveraged machine learning, specifically random forest classifiers, to automate event detection with accuracy comparable to manual coding [zemblys]. Additionally, methods like hidden Markov models (HMM) and minimum spanning trees offer nuanced classification by modeling temporal and spatial gaze patterns [10.1145/355017.355028]. Despite these innovations, traditional algorithms remain prevalent due to their simplicity and versatility, necessitating rigorous evaluation to optimize their performance across varied tasks [10.1145/355017.355028]. Recent comprehensive evaluations have demonstrated that algorithm choice significantly impacts classification outcomes [Andersson2017OneAlgorithm], motivating the need for systematic comparison studies.
This study evaluates the efficacy of I-VT, I-AVT, and I-DT for classifying fixations and saccades in eye-tracking data collected from free viewing and visual search tasks, and introduces an adaptive method for threshold selection based on minimization of state transitions, using a quantity we term the $K$-ratio. We then perform a Gaussian noise analysis to evaluate robustness (e.g., detecting performance degradation at noise levels $\geq$ 2 pixels).The analysis focuses on implementation difficulty, computational complexity, usability, and classification effectiveness, addressing the question: “Which algorithms are best suited for classifying eye-tracking features and movements?” Our findings aim to provide actionable insights for improving classification accuracy, advancing applications in cognitive research, clinical diagnostics, and assistive technologies.
@@ 2 Data and methods
Two paradigms were used for the same 15 participants. In a free-viewing condition (Random Pixel), participants viewed unstructured pixel fields (3 trials, 60 s each; $\sim$180,000 samples per participant, $\sim$2.7 M total in this subset). In a visual search condition (Waldo), participants located targets in complex scenes from Where’s Waldo? (9 trials, 45 s each; $\sim$405,000 samples per participant, $\sim$6 M total in this subset) [bookwaldo]. Each dataset comprised ten columns per sample at 1000 Hz: timestamp; left and right $x,y$ gaze in pixels; left and right pupil size; and event flags. See Tab. 1.
@@
The EyeLink software uses a combination of velocity and acceleration thresholds to classify saccades and fixations. We used the default settings (in degrees of visual angle: 30/s velocity, 8000/sec2 acceleration) to extract ground truth labels per frame.
To study the effect of noise in the classification algorithm, we augmented our data sample by generating a superposition of the collected trajectories with a tunable level of noise. These “noisy” trajectories were then analyzed additionally to the collected eye-gaze trajectories. We restricted to the case of Gaussian noise, with $\varepsilon_{x,i}$ and $\varepsilon_{y,i}$, added to each observation $i$, at the location $(x_{i},y_{i})$:
@@
where $\varepsilon_{x,i},\varepsilon_{y,i}\sim\mathcal{N}(0,\sigma^{2})$ are normally distributed with zero mean and a tunable standard deviation $\sigma$.
@@ 2.2 Classification algorithms
We introduce adaptive versions of the three classification algorithms mentioned above, I-VT, I-AVT, and I-DT [10.1145/355017.355028]. The I-VT algorithm uses the instantaneous velocity $V_{i}$, which is the velocity calculated between successive measurements, namely:
@@
where $x_{i},y_{i}$ are coordinates at timestamp $t_{i}$.
@@
The I-DT algorithm identifies fixations based on the spatial dispersion of gaze points within a moving time window. It uses two parameters, a minimum time window span $T_{\mathrm{min}}$ and a dispersion threshold $D_{max}$. The algorithm starts at the beginning of the time series and searches for the first index $i$ where
@@ 2.3 Optimization through a statistical metric
In this section, we introduce a metric we call $K$-ratio, which compares the empirical number of fixation-saccade transitions observed in an eye-gaze trajectory with the number of transitions between saccades and fixations that would be expected if data points simply alternated randomly between being fixations and saccades:
$\text{$K$-ratio}=\frac{p_{\mathrm{empirical}_{F\to S}}}{p_{\mathrm{independentF\rightarrow S}}},$ (7)
Figure 1: Example of scanpaths with overlaid classifications (orange = fixation, blue = saccade). Top row: Random Pixel task; bottom row: Waldo task. Columns (left to right): raw gaze points, I-VT, I-AVT, and I-DT results.
@@
where $n_{S}=N_{S}/N$, with $N_{S}$ the number of occurrences of the label $S$ and $N$ the total number of observations.
However, if the probability of transition from one state to another depends on the current state, this separation is not possible. The $K$-ratio is then
@@
where $n_{F\rightarrow S}$ is the number of fixation-to-saccade transitions divided by the number of data points.
By minimizing this ratio, thresholds used to label fixations and saccades are optimized towards keeping these two states separated in time.
@@
Figure 2: $K$-ratio curves used for threshold optimization (top: Random Pixel; bottom: Waldo). Minima define the task-specific thresholds for I-VT, I-AVT, and I-DT. The respective $K$-ratio for each of the algorithms for both task are tabulated in a supplementary section at the end of the paper. See Appendix Tab. S2. The $T_{\mathrm{min}}$ value used for I-DT is 50ms.
@@ 3 Results
Figure 3: Performance metrics at zero noise (left: Random Pixel; right: Waldo). Bars display overall accuracy, class-wise precision/recall/F1, and predicted fixation proportion for each algorithm.
@@
We evaluated I-VT, I-AVT, and I-DT on the Random Pixel (free viewing) and Waldo (visual search) datasets recorded at 1000 Hz, using the EyeLink event labels as frame-wise ground truth. The three classifiers were applied to the same recordings and compared in terms of threshold optimization, frame-wise accuracy, class-wise metrics, and robustness to added Gaussian noise. Frame-wise ground truth was provided by the standard EyeLink parser (30/s velocity; 8000/s2 acceleration).
@@
Figure 1 shows representative scanpaths with overlaid classifications for both tasks. I-VT produces compact fixation clusters with clearly delineated saccade segments. I-AVT tends to generate long fixation runs with relatively few saccade segments, while I-DT yields the most fixation-dominated segmentation, often merging brief saccades into extended fixation episodes. These qualitative patterns reflect the underlying class distributions and confusion trends that appear in the aggregate metrics. Thresholds were selected by minimizing the $K$-ratio for each algorithm and task. The $K$-ratio curves for the Random Pixel and Waldo datasets are shown in Fig. 2, and the corresponding optimal thresholds and $K$-ratio values are summarized in Appendix Tab. S2. In both tasks, the $K$-ratio decreases sharply at low thresholds and then levels off into a broad minimum. The Waldo task exhibits wider minima than the Random Pixel task, consistent with more heterogeneous oculomotor dynamics during structured visual search.
@@
Overall frame-wise accuracy and class-wise precision, recall, and F1-scores for fixations and saccades are summarized visually in Fig. 3. All algorithms slightly over-classify fixations compared to the ground truth proportion of 92%. Across both tasks, I-VT attains the highest overall accuracy and preserves near-complete recovery of fixations, while providing the strongest compromise between saccade precision and recall. I-AVT maintains high fixation performance but under-detects saccades, leading to low saccade F1-scores despite its use of directional information. I-DT shows a pronounced fixation bias, with the highest fixation proportions and low saccade recall, reflecting the tendency of the dispersion criterion to group short saccades into longer fixation segments. Taken together, the results indicate that I-VT offers the most balanced performance among the three methods when evaluated against the EyeLink labels.
@@
Figure 4: Robustness analysis of fixation classification algorithms under Gaussian spatial noise ($\sigma=0$-5050 pixels). Top row: Evolution of adaptive thresholds for I-VT, I-AVT, and I-DT algorithms across noise levels for Random Pixel and Waldo tasks. Bottom row: Classification accuracy comparing fixed thresholds (optimized at $\sigma=0$) versus adaptive thresholds (re-optimized at each noise level). Solid lines represent original (fixed) thresholds; dashed lines represent adaptive thresholds.
The effect of added Gaussian noise ($\sigma$ = 0-50 pixels) was examined both with the original threshold (rapid degradation) – cf. Tabs. S3-S6 in the supplementary section – and with thresholds re-optimized at each noise level via the $K$-ratio. See Fig. 4.
@@
Under clean recording conditions, I-VT yields the strongest agreement with the reference parser. When spatial noise is present, I-VT and I-AVT showed a progressive drift toward saccade-dominated classifications as noise increased, driven by inflated velocity and angular-velocity estimates. I-DT preserved more stable fixation proportions at intermediate noise levels but also deteriorated at higher amplitudes. With adaptive thresholding, I-DT shows the greatest stability in event proportions, while velocity-based algorithms remain markedly more sensitive to recording precision. However, in all cases, the $K$-ratio-based threshold at least partially mitigated against noise.

### v3

@@ Abstract
Properties of ocular fixations and saccades are highly stochastic during many experimental tasks, and their statistics are often used as proxies for various aspects of cognition. Although distinguishing saccades from fixations is not trivial, experimentalists generally use common ad-hoc thresholds in detection algorithms. This neglects inter-task and inter-individual variability in oculomotor dynamics, and potentially biases the resulting statistics. In this article, we introduce and evaluate an adaptive method based on a Markovian approximation of eye-gaze dynamics, using saccades and fixations as states such that the optimal threshold minimizes state transitions. Applying this to three common threshold-based algorithms (velocity, angular velocity and dispersion), we evaluate the overall accuracy against a multi-threshold benchmark as well as robustness to noise. We find that a velocity threshold achieves the highest baseline accuracy (90-93%) across both free-viewing and visual search tasks. However, velocity-based methods degrade rapidly under noise when thresholds remain fixed, with accuracy falling below 20% at high noise levels. Adaptive threshold optimization via $K$-ratio minimization substantially improves performance under noisy conditions for all algorithms. Adaptive dispersion thresholds demonstrate superior noise robustness, maintaining accuracy above 81% even at extreme noise levels ($\sigma=50$ px), though a precision-recall trade-off emerges that favors fixation detection at the expense of saccade identification. In addition to demonstrating our parsimonious adaptive thresholding method, these findings provide practical guidance for selecting and tuning classification algorithms based on data quality and analytical priorities.
@@ 1 Introduction
Conventional algorithms, such as Velocity-Threshold Identification (I-VT) and Dispersion-Threshold Identification (I-DT), rely on fixed velocity (e.g., 20–40°/s for I-VT) or spatial dispersion thresholds to classify eye movements [10.1145/355017.355028]. While computationally efficient and suitable for real-time applications, these methods are sensitive to noise and often fail to detect auxiliary events like microsaccades or post-saccadic oscillations (glissades), limiting their robustness in complex datasets [martinezs, nystrom2010adaptive]. Recent advancements have introduced more sophisticated approaches to address these limitations. For instance, Larsson et al. developed robust segmentation techniques for fixations and smooth pursuits in high-speed eye-tracking data, outperforming velocity- and dispersion-based methods [LARSSON2015145]. Zemblys et al. leveraged machine learning, specifically random forest classifiers, to automate event detection with accuracy comparable to manual coding [zemblys]. Additionally, methods like hidden Markov models (HMM) and minimum spanning trees offer nuanced classification by modeling temporal and spatial gaze patterns [10.1145/355017.355028]. Despite these innovations, traditional algorithms remain prevalent due to their simplicity and versatility, necessitating rigorous evaluation to optimize their performance across varied tasks [10.1145/355017.355028]. Recent comprehensive evaluations have demonstrated that algorithm choice significantly impacts classification outcomes [Andersson2017OneAlgorithm], motivating the need for systematic comparison studies.
This study evaluates the efficacy of threshold-based algorithms for classifying fixations and saccades in eye-tracking data collected from free viewing and visual search tasks, and introduces an adaptive method for threshold selection based on minimization of state transitions, using a quantity we term the $K$-ratio. We then perform a Gaussian noise analysis to evaluate robustness (e.g., detecting performance degradation at noise levels $\geq$ 2 pixels). The analysis focuses on implementation difficulty, computational complexity, usability, and classification effectiveness, addressing the question: “Which algorithms are best suited for classifying eye-tracking features and movements?” Our findings aim to provide actionable insights for improving classification accuracy, advancing applications in cognitive research, clinical diagnostics, and assistive technologies.
@@ 2 Data and methods
Eye movements were recorded at Oslo Metropolitan University in 2023–2024 under a Virtual-Eye project data-sharing agreement [rujeena2025-eye-gaze-dynamics-datasets]. More than 250 adults participated in the broader study; for the present report, we analyzed a matched within-subject subset of 15 participants who completed both tasks with high-quality data. The broader dataset, which includes the present tasks used in this study together with several additional experimental paradigms, has been published online as a data descriptor in Scientific Data [rujeena2025-eye-gaze-dynamics-datasets].
@@
Two paradigms were used for the same 15 participants. In a free-viewing condition (Random Pixel), participants viewed unstructured pixel fields (3 trials, 60 s each; $\sim$180,000 samples per participant, $\sim$2.7 M total in this subset). In a visual search condition (Waldo), participants located targets in complex scenes from Where’s Waldo? (9 trials, 45 s each; $\sim$405,000 samples per participant, $\sim$6 M total in this subset) [bookwaldo]. Each dataset comprised ten columns per sample at 1000 Hz: timestamp; left and right $x,y$ gaze in pixels; left and right pupil size; and event flags. Blinks and tracking loss are handled by removing samples with missing gaze coordinates (NaNs) before all feature computations. See Table 1.
@@
The EyeLink software uses a combination of velocity and acceleration thresholds to classify saccades and fixations. We used the default EyeLink parser settings (30/s velocity, 8000/sec2 acceleration) to extract reference event labels per frame, which serve as a practical benchmark rather than a physiological ground truth. To study the effect of noise in the classification algorithm, we augmented our data sample by generating a superposition of the collected trajectories with a tunable level of noise. These “noisy” trajectories were then analyzed additionally to the collected eye-gaze trajectories. We restricted to the case of Gaussian noise, with $\varepsilon_{x,i}$ and $\varepsilon_{y,i}$, added to each observation $i$, at the location $(x_{i},y_{i})$:
@@
where $\varepsilon_{x,i},\varepsilon_{y,i}\sim\mathcal{N}(0,\sigma^{2})$ are normally distributed with zero mean and a tunable standard deviation $\sigma$. $\sigma$ should be interpreted primarily as a controlled stress-test parameter for algorithmic robustness rather than a calibrated estimate of typical experimental measurement noise, although the smallest values $(\sigma\approx 1$–2 px) are in the range of plausible high-quality recordings.
@@ 2.2 Classification algorithms
We introduce adaptive versions of the three classification algorithms mentioned above, I-VT and I-DT [10.1145/355017.355028], as well as an angle-weighted variant of I-VT which we call Angular Velocity-Threshold Identification (I-AVT). The I-VT algorithm uses the instantaneous velocity $V_{i}$, which is the velocity calculated between successive measurements, namely:
@@
where $x_{i},y_{i}$ are coordinates at timestamp $t_{i}$. All kinematic features used by the threshold algorithms are computed from the recorded screen coordinates and are therefore expressed in pixel-based units (px/s for velocity-like quantities), unless explicitly converted to degrees of visual angle.
@@
Savitzky-Golay smoothing filter was applied before computing I-AVT to reduce high-frequency fluctuations at 1000 Hz. The cosine term is computed only for nonzero displacement vectors; samples with zero-length steps are excluded, and the effective velocity is taken.
The I-DT algorithm identifies fixations based on the spatial dispersion of gaze points within a moving time window. It uses two parameters, a minimum time window span $T_{\mathrm{min}}$ and a dispersion threshold $D_{max}$. We set $T_{\mathrm{min}}=50\,\mathrm{ms}$, which corresponds to 50 samples at 1000 Hz and serves to suppress spurious ‘fixations’ caused by brief noise-driven clustering or microsaccade-scale interruptions. This value is also well below typical fixation durations (hundreds of ms), so it acts as a conservative lower bound rather than a tuning knob that shapes the main results. The algorithm starts at the beginning of the time series and searches for the first index $i$ where
@@ 2.3 Optimization through a statistical metric
In this section we introduce a metric we call $K$-ratio, which compares the empirical number of fixation-saccade transitions observed in an eye-gaze trajectory with the number of transitions between saccades and fixations that would be expected if data points simply alternated randomly between being fixations and saccades:
$\text{$K$-ratio}=\frac{p_{\mathrm{empirical}_{F\to S}}}{p_{\mathrm{independentF\rightarrow S}}}.$ (7)
Figure 1: Example of scanpaths with overlaid classifications (orange = fixation, blue = saccade). Top row: Random Pixel task; bottom row: Waldo task. Columns (left to right): raw gaze points, I-VT, I-AVT, and I-DT results. All gaze positions are in screen pixels.
@@
where $n_{S}=N_{S}/N$, with $N_{S}$ the number of occurrences of the label $S$ and $N$ the total number of observations. However, if the probability of transition from one state to another depends on the current state, this separation is not possible. The $K$-ratio is then
@@
where $n_{F\rightarrow S}$ is the number of fixation-to-saccade transitions divided by the number of data points. We therefore assessed time-local stability of transition statistics and the short-lag adequacy of a first-order Markov proxy by repeating the transition computations in equal-duration blocks and by comparing empirical $k$-step transitions to the Markov prediction $T^{k}$ (see Appendix A, Fig. A1). By minimizing this ratio, thresholds used to label fixations and saccades are selected towards keeping these two states separated in time. This approach involves a methodological trade-off: it enhances coherence and suppresses noise-induced artifacts at the expense of sensitivity to high-frequency micro-movements or post-saccadic oscillations. In noisy recordings, this can be beneficial as a form of regularization, but it also implies a trade-off in which short or noise-contaminated saccades may be absorbed into neighboring fixations.
@@
Figure 2: $K$-ratio curves used for threshold optimization (top: Random Pixel; bottom: Waldo). Minima define the task-specific thresholds for I-VT, I-AVT, and I-DT. The respective $K$-ratio for each of the algorithms for both task are tabulated in a supplementary section at the end of the paper. See Appendix Table 3. For I-DT we use $T_{\mathrm{min}}=50\,\mathrm{ms}$. Threshold units are in pixels for I-AVT, I-VT (px/s), and I-DT (px).
@@ 3 Results
Figure 3: Performance metrics at zero noise (left: Random Pixel; right: Waldo) using EyeLink reference labels. Bars display overall accuracy, class-wise precision/recall/F1, and predicted fixation proportion for each algorithm. Since fixation dominates ($\sim 92\%$), the saccade F1-score ($\mathrm{F1}$) and its precision/recall components are the most informative summary of rare-event detection performance.
@@
We evaluated I-VT, I-AVT, and I-DT on the Random Pixel (free viewing) and Waldo (visual search) datasets recorded at 1000 Hz, using the EyeLink event labels as frame-wise reference labels. The three classifiers were applied to the same recordings and compared in terms of threshold optimization, frame-wise accuracy, class-wise metrics, and robustness to added Gaussian noise. Frame-wise reference labels were provided by the standard EyeLink parser (30/s velocity; 8000/s2 acceleration).
@@
Figure 1 shows representative scanpaths with overlaid classifications for both tasks. I-VT produces compact fixation clusters with clearly delineated saccade segments. I-AVT tends to generate long fixation runs with relatively few saccade segments, while I-DT yields the most fixation-dominated segmentation, often merging brief saccades into extended fixation episodes. These qualitative patterns reflect the underlying class distributions and confusion trends that appear in the aggregate metrics. Thresholds were selected by minimizing the $K$-ratio for each algorithm and task. The $K$-ratio curves for the Random Pixel and Waldo datasets are shown in Fig. 2, and the corresponding optimal thresholds and $K$-ratio values are summarized in Appendix Table 3. In both tasks, the $K$-ratio decreases sharply at low thresholds and then levels off into a broad minimum. The Waldo task exhibits wider minima than the Random Pixel task, consistent with more heterogeneous oculomotor dynamics during structured visual search.
@@
Overall frame-wise accuracy and class-wise precision, recall, and F1-scores for fixations and saccades are summarized visually in Fig. 3. Because the datasets are strongly fixation-dominated ($\sim$92%), overall accuracy is driven primarily by fixation performance and is therefore secondary for model comparison. We treat the saccade F1-score ($\mathrm{F1}_{S}$) as the primary metric, and interpret accuracy only as a contextual summary. Under noise-free conditions, I-VT achieves the highest $\mathrm{F1}_{S}$ among the three methods (Random Pixel: $\mathrm{F1}_{S}=0.51$; Waldo: $\mathrm{F1}_{S}=0.54$), compared with I-AVT (0.46, 0.53) and I-DT (0.42, 0.48) (Table 5 - 7). Under increasing noise, all methods show declining $\mathrm{F1}_{S}$, with adaptive thresholding partially mitigating the degradation, but with a pronounced trade-off for I-DT where fixation performance remains high while saccade detection becomes conservative at high $\sigma$ (Table 7 and Fig. 9). All algorithms slightly over-classify fixations compared to the reference label proportion of 92%. Across both tasks, I-VT generally showed strong rare-event performance while maintaining high fixation performance. I-AVT maintains high fixation performance but under-detects saccades, leading to low saccade F1-scores despite its use of directional information. I-DT shows a pronounced fixation bias, with the highest fixation proportions and low saccade recall, reflecting the tendency of the dispersion criterion to group short saccades into longer fixation segments. Taken together, the results indicate that I-VT provides a favorable overall balance among the three methods when evaluated against the EyeLink labels, although the relative differences between methods depend on task and noise condition. Both the optimal thresholds and obtained $K$-ratio minima showed significant inter-participant differences. The adaptive I-VT algorithm in particular showed consistently high intraclass correlations, with ICC(1,1) values of 0.688 and 0.728 for the Random Pixel and Waldo task thresholds, respectively, and 0.816 and 0.853, respectively, for their $K$-ratios (see Appendix B).
@@
Figure 4: Robustness analysis of fixation classification algorithms under Gaussian spatial noise ($\sigma=0$-5050 pixels). Top row: Evolution of adaptive thresholds for I-VT, I-AVT, and I-DT algorithms across noise levels for Random Pixel and Waldo tasks. Bottom row: Classification accuracy comparing fixed thresholds (optimized at $\sigma=0$) versus adaptive thresholds (re-optimized at each noise level). Solid lines represent original (fixed) thresholds; dashed lines represent adaptive thresholds. Noise amplitude $\sigma$ is specified in screen pixels (px), since perturbations are applied directly to the recorded (x,y) pixel coordinates.
@@
The effect of added Gaussian noise ($\sigma$ = 0-50 pixels) was examined both with the original threshold (rapid degradation) – cf. Tabs. 4-7 in the supplementary section – and with thresholds re-optimized at each noise level via the $K$-ratio. See Fig. 4. Under clean recording conditions, I-VT yields the highest agreement with the reference parser. When spatial noise is present, I-VT and I-AVT showed a progressive drift toward saccade-dominated classifications as noise increased, driven by inflated velocity and angular-velocity estimates. I-DT preserved more stable fixation proportions at intermediate noise levels but also deteriorated at higher amplitudes. To quantify the corresponding change in event structure, we measured fixation-duration distributions under noise for I-VT with fixed thresholds the median fixation duration collapses from 619 ms at $\sigma=0$ to 1 ms for $\sigma\geq 5$ (with $\Pr[T_{\mathrm{fix}}\geq 200~\mathrm{ms}]=0$ from $\sigma\geq 2$), whereas adaptive I-DT retains substantially longer fixations, with median 318 ms at $\sigma=0$ and 431 ms at $\sigma=50$ and $\Pr[T_{\mathrm{fix}}\geq 200~\mathrm{ms}]$ increasing from 0.887 ($\sigma=0$) to 0.929 ($\sigma=50$), consistent with fixation stretching via saccade absorption. With adaptive thresholding, I-DT shows the greatest stability in event proportions, while velocity-based algorithms remain markedly more sensitive to recording precision. However, in all cases, the $K$-ratio-based threshold at least partially mitigated against noise.